1. Let f(x) = 2x^2 − 8x − 7.
What is the quadratic in vertex form that is y = a(x − h)^2 + k
for some constants a, h, and k). Use the vertex form to locate the x intercepts of the function. (without the quadratic formula.) Sketch the graph of f, showing all important features. Solve f(x) > 0.

Answer by josgarithmetic(39617)

Learn to change y=ax^2+bx+c general form into y=a(x-h)^2+k standard form.

This shows how to make the change.

The vertex is the point (h,k) and find the x-intercepts from the standard form by setting y=9 and solve for x.